OpenStudy (anonymous):

what is the y intercept of 2y-5x=14

OpenStudy (anonymous):

hi welcome to openstudy @gcam11 :) Find the y-int by substituting x=0 in the equation and solve for y.

OpenStudy (anonymous):

It will be like this: \(\sf 2y-5x=14\\2y-5(0)=14\) So what will you get when x=0?

OpenStudy (anonymous):

2y-0=14

OpenStudy (anonymous):

2y=14 now evaluate y by dividing both sides by 2

OpenStudy (anonymous):

1y=7

OpenStudy (anonymous):

yes, so the y int is 7 or point (0,7) :)

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

you're welcome :)

OpenStudy (anonymous):

would send u a metal but dont know how too